# Diff-PCR: Diffusion-Based Correspondence Searching in Doubly Stochastic Matrix Space for Point Cloud Registration

### v4

@@ III-D Diffusion Model in Doubly Stochastic Matrix space
$&logp(E^{0})\geq\mathbb{E}_{E^{1:T}\sim q(E^{1:T}|E^{0})}\left[log\left(\frac{p_{\theta}(E^{0:T})}{q(E^{1:T}|E^{0})}\right)\right]\\ &=\mathbb{E}_{q}\left[log\left(\frac{p_{\theta}(E^{0:T})}{q(E^{1:T}|E^{0})}\right)\right]\\ &=\mathbb{E}_{q}\left[logp_{\theta}(E^{T})+log(\frac{p_{\theta}(E^{0}|E^{1})}{q(E^{1}|E^{0})})+\sum\limits_{t=2}^{T}log(\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\theta}(E^{t-1}|E^{t})}}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}q(E^{t}|E^{t-1})}})\right]\\ &=\mathbb{E}_{q}\left[logp_{\theta}(E^{T})+log(\frac{p_{\theta}(E^{0}|E^{1})}{q(E^{1}|E^{0})})\right]\\ &+\mathbb{E}_{q}\left[\sum\limits_{t=2}^{T}log\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\frac{q(E^{t-1}|E^{0})}{q(E^{t}|E^{0})}}\cdot\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}q(E^{t-1}|E^{t},E^{0})p_{\theta}(E^{0}|E^{t})}}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}q(E^{t-1}|E^{t},E^{0})}}\right)\right]\\ &=\mathbb{E}_{q}\left[S_{\theta}(E^{T},E^{1},E^{0})+\sum\limits_{t=2}^{T}log(p_{\theta}(E^{0}|E^{t}))\right]\end{aligned}$

### v5

@@ III-D Diffusion Model in Doubly Stochastic Matrix space
$&logp(E^{0})\geq\mathbb{E}_{E^{1:T}\sim q(E^{1:T}|E^{0})}\left[log\left(\frac{p_{\theta}(E^{0:T})}{q(E^{1:T}|E^{0})}\right)\right]\\ &=\mathbb{E}_{q}\left[log\left(\frac{p_{\theta}(E^{0:T})}{q(E^{1:T}|E^{0})}\right)\right]\\ &=\mathbb{E}_{q}\left[logp_{\theta}(E^{T})+log(\frac{p_{\theta}(E^{0}|E^{1})}{q(E^{1}|E^{0})})+\sum\limits_{t=2}^{T}log(\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p_{\theta}(E^{t-1}|E^{t})}}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}q(E^{t}|E^{t-1})}})\right]\\ &=\mathbb{E}_{q}\left[logp_{\theta}(E^{T})+log(\frac{p_{\theta}(E^{0}|E^{1})}{q(E^{1}|E^{0})})\right]\\ &+\mathbb{E}_{q}\left[\sum\limits_{t=2}^{T}log\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{q(E^{t-1}|E^{0})}{q(E^{t}|E^{0})}}\cdot\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}q(E^{t-1}|E^{t},E^{0})p_{\theta}(E^{0}|E^{t})}}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}q(E^{t-1}|E^{t},E^{0})}}\right)\right]\\ &=\mathbb{E}_{q}\left[S_{\theta}(E^{T},E^{1},E^{0})+\sum\limits_{t=2}^{T}log(p_{\theta}(E^{0}|E^{t}))\right]\end{aligned}$